A sample of 500 field mice contains 245 individuals that are D1D1, 165 that are D1D2, and 90 that are D2D2.

A. What is the frequency of D1 in this sample? Express your answer using three decimal places.

B. What is the frequency of D2 in this sample? Express your answer using three decimal places.

C. Is this population in Hardy-Weinberg equilibrium? Use the chi-square test to justify your answer.

D. Is inbreeding a possible genetic explanation for the observed distribution of genotypes?
Answer #1

Please look in the image for solution:

A • - frequency of Di = genotypic frequency of 0,0, + +0,02 = 24,5 + Ź x oo 500 2 = 0.49 +4x 0.33 = 0.49 +0.165 - 0-655 Allel